OpenStudy (1645323):

Factor out x^3-27x^6

OpenStudy (1645323):

@hartnn @Callisto @phi

hartnn (hartnn):

first factor out x^3 from both the terms, what u get ?

OpenStudy (1645323):

actuall thats the wrong equation: its 8x^3-27x^6

hartnn (hartnn):

still, you'll be able to factor out x^3

OpenStudy (1645323):

x^3(8-27x^3)

hartnn (hartnn):

yes, 8 is the cube of which number ? 27 is the cube of which number x is the cube of which variable ?

OpenStudy (1645323):

would it be X^3(2-3x)(4+6x+9x^2) and how would I solve this to find the solutions?

hartnn (hartnn):

thats correct, and why do u need to solve it ? when the question just says to factorize ?

OpenStudy (1645323):

Thats what my teacher wants to to do, she said find the solution

hartnn (hartnn):

can you solve quadratics ? know the quadratic formula ?

OpenStudy (1645323):

yes so I just solve it using the new equation or original equation?

hartnn (hartnn):

\( x^3(2-3x)(4+6x+9x^2) = 0 \\ x^3 = 0, \quad x=...? \\ 2-3x= 0, \quad x=...? \\(4+6x+9x^2) = 0, use ~ quadratic formula \)

OpenStudy (1645323):

ok thanks

hartnn (hartnn):

welcome ^_^